OpenStudy (anonymous):

Steve purchased $61.50 worth of supplies for a hiking trip. The supplies included flashlights for $3.50 each, compasses for $1.50 each, and water bottles for $0.75 each. a. Write a linear equation in three variables to represent this situation. b. Steve purchased 6 flashlights and 24 water bottles. How many compasses did he purchase?

OpenStudy (anonymous):

Part A: You know he spent $61.50 and the price for each. Let F be the number of flashlights, C be the number of compasses, and W be the number of water bottles. 3.5F + 1.5C + 0.75W = 61.50

OpenStudy (anonymous):

Part B: Substitute 6 for F and 24 for W and solve for C to get the number of compasses he bought. 3.5(6) + 1.5C + 0.75(24) = 61.50 21 + 1.5C + 18 = 61.50 1.5C = 22.5 C = 15